# Pulsar Timing Arrays require hierarchical models

### v1

@@ 4.2 First IPTA Mock Data Challenge
where the last two lines give the two different options for spin noise amplitude: either log-uniform priors, or uniform priors. The parameters $log_{10}n_{a}$ represent the spin noise amplitude of pulsar $a$ in the same units that Enterprise uses. In those units, the spin noise is basically defined as a pulsar-term GW signal for only pulsar $a$. For details on how the likelihood and posterior are formed for realistic data, I refer to van Haasteren & Vallisneri (2014, and references therein).
@@
From the MDC1 onward, the log-uniform priors were widely accepted as the “correct” choice for spin noise and many other noise processes for detection, with the acknowledgement that these priors are effectively improper and need arbitrary bounds. When upper-limits on a parameter need to be set, current software packages allow somewhat arbitrarily to switch to uniform priors so that the lower bound of amplitude parameters is well-defined. In Figure 4 I have also plotted the posterior distribution that is obtained with a model without any spin noise, labeled as “Physical”. The full posterior here is only one-dimensional, meaning that no marginalization or sampling was required to make it. It is clear that, although the “Log-Uniform” and “Physical” priors give comparable results, the posterior made with the “Physical” prior is shifted slightly to the right.
The hierarchical model I used for the MDC was:
@@
This example demonstrates that both uninformative priors that were applied to the IPTA MDC1 datasets in the literature were sub-optimal in hindsight. It just so happened that the Log-Uniform priors were not wrong enough for anyone to notice until now. Using nested sampling with Dynesty (Speagle, 2020), I have determined that the Bayes factor $\mathcal{B}=Z_{\rm{Hierarchical}}/Z_{\rm{Log-Uniform}}\approx 1665$ is significantly in favor of the Hierarchical model. Note that, since MDC1 did not contain any injected spin noise, the likelihood is effectively “high” in a large portion of prior space: everywhere under the $\log_{10}n_{a}\leq-13.7$ the likelihood value is somewhat similar, since that is small with respect to the injected GWB. If spin noise were present and detectable, like it is in realistic datasets, the Bayes factors between models with Hierarchical and Log-Uniform priors can get drastically larger.
@@ 4.3 More realistic mock data
As a last more realistic test, I generated some mock data that was loosely modeled after the NANOGrav 12.5yr dataset (Alam et al., 2020), which showed a significant detection of a common process in the GWB search (Arzoumanian et al., 2020). Correlations were not significantly observed, but a model with a common uncorrelated process (CURN) was preferred over noise-only with a Bayes factor $\mathcal{B}=Z_{\rm{CURN}}/Z_{\rm{Noise}}>10^{4}$. I created a custom Enterprise class for mock data so that no actual PTA datasets were required as input. As timing model I used quadratic spindown and astrometry parameters, the latter of which were created using Astropy (Astropy Collaboration et al., 2022) [^1] Code to generate mock data as an Enterprise addition:
@@
. Time of observations and spin noise parameters were loosely based on the NANOGrav 12.5yr datasets. Spin noise parameters $\log_{10}A$ and $\gamma$ are typically covariant due to the choice of reference frequency $f_{r}$. I generated the spin noise parameters loosely scattered around the line of covariance with total timing residual variance based on the values in the literature. The exact values are shown in Figure 5. For the GWB, I injected a $\log_{10}h_{c}=-15$ signal that is correlated among pulsars with the regular Hellings & Downs correlations $\Gamma(\zeta)$.
@@
Using a different random number seed at the start of the data generation, I created several realizations to see how often I would get a discrepancy between a model with Log-Uniform priors and Hierarchical priors. It turned out that occasionally both models were in agreement, and sometimes they did not. The toy model analysis from Section 2.1, which was built completely independently with PyMC, exhibited the same behavior. For the mock data in this Section, I noticed a serious discrepancy between the model with Log-Uniform priors and the model with Hierarchical priors on the third realization of mock data, which gives an impression of how often such posterior bias occurs. The analysis result of that mock dataset is shown in Figure 6. It is clear that only the model with Hierarchical priors managed to recover the injection here. Statistical analysis coupled with more proper noise modeling of realistic PTA data is postponed to future investigations. The goal of this example is to strengthen the statement made in the title of this letter: posterior bias is not rare, and an HBM can address it.
@@
To show the effect of the HBM on the recovery of spin noise parameters I constructed point estimators of $\log_{10}A$ and $\gamma$ for each pulsar under the different priors, as can be seen in Figure 5. For the Log-Uniform priors, I used the Maximum Likelihood Estimator (MLE, marked with ’o’) on the likelihood from Enterprise. For the HBM (marked with ’x’), I used the log-posterior distribution minus the log-Jacobian to avoid artifacts coming from the coordinate transformation around the edges of the allowed interval for the $\gamma$ parameters. I also plotted the injected values (marked with ’+’). The effect of shrinkage can be seen, where the HBM estimates in general are more clustered towards the middle of the distribution. One caveat with these results is that the CURN model is lightly degenerate with respect to the model parameters, so the optimization results are not very stable numerically. Some point estimates have a significant margin of error. Regardless, the shrinkage effect can be observed.
@@
Figure 5: Point estimators for the spin noise parameters $(\log_{10}A,\gamma)$ for the Log-Uniform (Maximum Likelihood Estimator, MLE) and the HBM model, plotted together with the injected values. The Posterior Predictive for $(\log_{10}A,\gamma)$ is shown as contours. Although the point estimates were not very stable numerically due to light degeneracies in the model, the effects of shrinkage can be observed as the HBM point estimates are drawn towards the middle of the posterior predictive. For some pulsars I have also drawn lines between the MLE, HBM, and injection values, so that their relationship can be seen.
@@ 5 Comparison with attempts in the literature
Several studies in the literature have attempted to explain potential biases in PTA results. Firstly, Zic et al. (2022) pointed out that spurious detections could happen in simulated datasets that did not have a common signal injected. Similarly, Johnson et al. (2022) point out that biases can happen when analyzing PTA data, and they evaluate many realizations of data to understand when this is happening. Goncharov et al. (2022) attempt to remedy biases found by Zic et al. (2022) by including a “quasi-common process”, which is similar to the common process (GWB) but with amplitudes governed by hyperparameters. Interestingly, they actually modeled the common process with a hierarchical model that is close to what they should have used for the noise (except that spectral index was linked). Lastly, Hazboun et al. (2020) found posterior bias in their simulations, and they discovered that a “dropout” model, that allows the data to decide whether the model should include a spin noise component for a pulsar, can reduce the posterior bias in the GW parameters. The approach of turning off the spin noise of a pulsar (as if removing it from the ensemble) based on the single-pulsar analysis odds ratio is incorrect. This is because this process is equivalent to using the data to change the prior odds of various noise models, and then re-analyzing said data with updated prior odds informed by the data. In the Bayesian Statistics literature having a data-informed prior is referred to as double-dipping, and it is not allowed. In general,
@@ 6 Implications for Pulsar Timing Array projects
The observed correlations reported by the various PTA projects, such as those calculated with the Optimal Statistic (Anholm et al., 2009; Chamberlin et al., 2015; Allen & Romano, 2023; Vigeland et al., 2018; Sardesai et al., 2023; Vallisneri et al., 2023), also depend on the noise parameters. However, even with biased noise parameter values, those statistics are still unbiased. And, the false alarm probabilities created using simulations based on the null-Hypothesis should still hold if the background estimation method from sky scrambles and phase shifts (Cornish & Sampson, 2016; Taylor et al., 2017) are valid. Although this has not been formally derived (Di Marco et al., 2023), simulations show that such background estimates are quite reasonable (Taylor, private communications). The concerned reader can therefore have some peace of mind that the evidence for a GWB in PTA data is not going to go away. Regardless, even the Noise Marginalized Optimal Statistic will have to be recalculated with updated noise parameters. It is unclear what results will change to when appropriate HBMs are crafted for real PTA data. This will be a community effort, and the model explored in this letter is almost surely inadequate in some way. Regardless, because the model has been misspecified when uninformative priors are used on parameters that are covariant with the GWB, the inference on GWB parameters will change under an HBM. At the end of the line, prepare for updated Bayes Factors, modified estimates of the characteristic strain $h_{c}$ for power-law models, updated GWB spectral estimates in general, and potentially slightly altered false alarm probabilities.
@@ 7 Conclusions
Gravitational-Wave searches by Pulsar Timing Arrays use an array of pulsars which are jointly analyzed. Many noise sources are modeled simultaneously with the gravitational wave signal. The noise source processes originating from individual pulsars are separately characterized in such an analysis, and the parameters of these processes are typically given an uninformative prior of some sort. Such uninformative priors should not be used when an ensemble of processes is modeled jointly, and instead the underlying distribution of noise parameters of the ensemble of pulsars should be modeled simultaneously with all the pulsar properties. This joint model can be elegantly described by a Hierarchical Bayesian Model, which is the combination of the original likelihood with a different Hierarchical Prior. Neglecting to model the underlying distribution in such a way biases model parameters. Especially Gravitational-Wave parameters are affected, since the GWB is somewhat covariant with the whole ensemble of spin noise parameters. Current results in the PTA literature all use uninformative priors for the pulsar spin noise, and should be re-evaluated in the context of hierarchical models. This could result in updated Bayes Factors, different (GWB) spectral estimates, and slightly modified false alarm probabilities.

### v2

@@ 4.2 First IPTA Mock Data Challenge
where the last two lines give the two different options for spin noise amplitude: either log-uniform priors, or uniform priors. The parameters $log_{10}n_{a}$ represent the spin noise amplitude of pulsar $a$ in the same units that Enterprise uses. In those units, the spin noise is basically defined as a pulsar-term GW signal for only pulsar $a$. For details on how the likelihood and posterior are formed for realistic data, we refer to van Haasteren & Vallisneri (2014, and references therein).
@@
From the MDC1 onward, the log-uniform priors were widely accepted as the “correct” choice for spin noise and many other noise processes for detection, with the acknowledgement that these priors are effectively improper and need arbitrary bounds. When upper-limits on a parameter need to be set, current software packages allow somewhat arbitrarily to switch to uniform priors so that the lower bound of amplitude parameters is well-defined. In Figure 4 we have also plotted the posterior distribution that is obtained with a model without any spin noise, labeled as “Physical”. The full posterior here is only one-dimensional, meaning that no marginalization or sampling was required to make it. It is clear that, although the “Log-Uniform” and “Physical” priors give comparable results, the posterior made with the “Physical” prior is shifted slightly to the right.
The hierarchical model we used for the MDC was:
@@
This example demonstrates that both uninformative priors that were applied to the IPTA MDC1 datasets in the literature were sub-optimal in hindsight. It just so happened that the Log-Uniform priors were not wrong enough for anyone to notice until now. Using nested sampling with Dynesty (Speagle, 2020), we have determined that the Bayes factor $\mathcal{B}=Z_{\rm{Hierarchical}}/Z_{\rm{Log-Uniform}}\approx 1665$ is significantly in favor of the Hierarchical model. Note that, since MDC1 did not contain any injected spin noise, the likelihood is effectively “high” in a large portion of prior space: everywhere under the $\log_{10}n_{a}\leq-13.7$ the likelihood value is somewhat similar, since that is small with respect to the injected GWB. If spin noise were present and detectable, like it is in realistic datasets, the Bayes factors between models with Hierarchical and Log-Uniform priors can get drastically larger.
@@ 4.3 More realistic mock data
As a last more realistic test, we generated some mock data that was loosely modeled after the NANOGrav 12.5yr dataset (Alam et al., 2020), which showed a significant detection of a common process in the GWB search (Arzoumanian et al., 2020). Correlations were not significantly observed, but a model with a common uncorrelated process (CURN) was preferred over noise-only with a Bayes factor $\mathcal{B}=Z_{\rm{CURN}}/Z_{\rm{Noise}}>10^{4}$. We created a custom Enterprise class for mock data so that no actual PTA datasets were required as input. As timing model we used quadratic spindown and astrometry parameters, the latter of which were created using Astropy (Astropy Collaboration et al., 2022) [^1] Code to generate mock data as an Enterprise addition:
@@
. Time of observations and spin noise parameters were loosely based on the NANOGrav 12.5yr datasets. Spin noise parameters $\log_{10}A$ and $\gamma$ are typically covariant due to the choice of reference frequency $f_{r}$. We generated the spin noise parameters loosely scattered around the line of covariance with total timing residual variance based on the values in the literature. The exact values are shown in Figure 5. For the GWB, we injected a $\log_{10}h_{c}=-15$ signal that is correlated among pulsars with the regular Hellings & Downs correlations $\Gamma(\zeta)$.
@@
Using a different random number seed at the start of the data generation, we created several realizations to see how often we would get a discrepancy between a model with Log-Uniform priors and Hierarchical priors. It turned out that occasionally both models were in agreement, and sometimes they did not. The toy model analysis from Section 2.1, which was built completely independently with PyMC, exhibited the same behavior. For the mock data in this Section, we noticed a serious discrepancy between the model with Log-Uniform priors and the model with Hierarchical priors on the third realization of mock data, which gives an impression of how often such posterior bias occurs. The analysis result of that mock dataset is shown in Figure 6. It is clear that only the model with Hierarchical priors managed to recover the injection here. Statistical analysis coupled with more proper noise modeling of realistic PTA data is postponed to future investigations. The goal of this example is to strengthen the statement made in the title of this letter: posterior bias is not rare, and an HBM can address it.
@@
To show the effect of the HBM on the recovery of spin noise parameters we constructed point estimators of $\log_{10}A$ and $\gamma$ for each pulsar under the different priors, as can be seen in Figure 5. For the Log-Uniform priors, we used the Maximum Likelihood Estimator (MLE, marked with ’o’) on the likelihood from Enterprise. For the HBM (marked with ’x’), we used the log-posterior distribution minus the log-Jacobian to avoid artifacts coming from the coordinate transformation around the edges of the allowed interval for the $\gamma$ parameters. We also plotted the injected values (marked with ’+’). The effect of shrinkage can be seen, where the HBM estimates in general are more clustered towards the middle of the distribution. One caveat with these results is that the CURN model is lightly degenerate with respect to the model parameters, so the optimization results are not very stable numerically. Some point estimates have a significant margin of error. Regardless, the shrinkage effect can be observed.
@@
Figure 5: Point estimators for the spin noise parameters $(\log_{10}A,\gamma)$ for the Log-Uniform (Maximum Likelihood Estimator, MLE) and the HBM model, plotted together with the injected values. The Posterior Predictive for $(\log_{10}A,\gamma)$ is shown as contours. Although the point estimates were not very stable numerically due to light degeneracies in the model, the effects of shrinkage can be observed as the HBM point estimates are drawn towards the middle of the posterior predictive. For some pulsars we have also drawn lines between the MLE, HBM, and injection values, so that their relationship can be seen.
@@ 5 Comparison with attempts in the literature
Several studies in the literature have attempted to explain potential biases in PTA results. Firstly, Zic et al. (2022) pointed out that spurious detections could happen in simulated datasets that did not have a common signal injected. Similarly, Johnson et al. (2022) point out that biases can happen when analyzing PTA data, and they evaluate many realizations of data to understand when this is happening. Goncharov et al. (2022) remedy the model misspecification found by Zic et al. (2022) by including a “quasi-common process”, which is similar to the common process (GWB) but with amplitudes governed by hyperparameters. Uninformative priors were left in place for noise. Such a model seemed to resolve the posterior bias they found, but it is still misspecified from a population point of view. Lastly, Hazboun et al. (2020) found posterior bias in their simulations, and they discovered that a “dropout” model, that allows the data to decide whether the model should include a spin noise component for a pulsar, can reduce the posterior bias in the GW parameters. The approach of turning off the spin noise of a pulsar (as if removing it from the ensemble) based on the single-pulsar analysis odds ratio is incorrect. This is because this process is equivalent to using the data to change the prior odds of various noise models, and then re-analyzing said data with updated prior odds informed by the data. In the Bayesian Statistics literature having a data-informed prior is referred to as double-dipping, and it is not allowed. In general,
@@ 6 Implications for Pulsar Timing Array projects
The observed correlations reported by the various PTA projects, such as those calculated with the Optimal Statistic (Anholm et al., 2009; Chamberlin et al., 2015; Allen & Romano, 2023; Vigeland et al., 2018; Sardesai et al., 2023; Vallisneri et al., 2023), also depend on the noise parameters. However, even with biased noise parameter values, those statistics are still unbiased. Further, the false alarm probabilities created using simulations based on the null-Hypothesis should still hold if the background estimation method from sky scrambles and phase shifts (Cornish & Sampson, 2016; Taylor et al., 2017) are valid. Although this has not been formally derived (Di Marco et al., 2023), simulations show that such background estimates are quite reasonable (Taylor, private communications). The concerned reader can therefore have some peace of mind that the evidence for a GWB in PTA data is not going to go away. Regardless, even the Noise Marginalized Optimal Statistic will have to be recalculated with updated noise parameters. It is unclear what results will change to when appropriate HBMs are crafted for real PTA data. This will be a community effort, and the model explored in this letter is almost surely inadequate in some way. Regardless, because the model has been misspecified when uninformative priors are used on parameters that are covariant with the GWB, the inference on GWB parameters will change under an HBM. In summary, one should prepare for updated Bayes Factors, modified estimates of the characteristic strain $h_{c}$ for power-law models, updated GWB spectral estimates in general, and potentially slightly altered false-alarm probabilities.
@@ 7 Conclusions
GW searches by PTAs use an array of pulsars which are jointly analyzed. Many noise sources are modeled simultaneously with the GW signal. The noise source processes originating from individual pulsars are separately characterized in such an analysis, and the parameters of these processes are typically given an uninformative prior of some sort. Such uninformative priors should not be used when an ensemble of processes is modeled jointly, and instead the underlying distribution of noise parameters of the ensemble of pulsars should be modeled simultaneously with all the pulsar properties. This joint model can be elegantly described by a Hierarchical Bayesian Model, which is the combination of the original likelihood with a different Hierarchical Prior. Neglecting to model the underlying distribution in such a way biases model parameters. Especially GW parameters are affected, since the GWB is somewhat covariant with the whole ensemble of spin noise parameters. Current results in the PTA literature all use uninformative priors for the pulsar spin noise, and should be re-evaluated in the context of hierarchical models. This could result in updated Bayes Factors, different (GWB) spectral estimates, and slightly modified false alarm probabilities.